illy rides his bicycle east on a bike path that is 75 kilometers long on a bike path but goes back 5 km, west, to sit on a bench he passed. If it takes Billy 15 hours to do so, then: 

 

a. What is his speed?  

 

 

 

b. What is his velocity?  

 


1
Expert's answer
2021-08-09T06:52:19-0400

a)


"V=\frac{75+5}{15}=5.3\frac{km}{h}"

b)


"v=\frac{75-5}{15}=4.7\frac{km}{h}"